OpenStudy (anonymous):

Am I right?! The position of an object at time t is given by s(t) = 1 - 12t. Find the instantaneous velocity at t = 2 by finding the derivative. I think the answer is 6.

OpenStudy (sleepyhead314):

what did you get for the derivative of 1-12t ?

OpenStudy (anonymous):

I assumed it would be sqrt(12) so I got 3.46

OpenStudy (kainui):

Why would you think to assume that @TheHero ?

OpenStudy (anonymous):

Because I know when I'm looking for the derivative of 8/x x=9 for example, it would be 8/9^2

OpenStudy (sleepyhead314):

that is true only because the derivative of 1/x = 1/x^2 meanwhile if you have -15x derivative of that is just -15 power rule

OpenStudy (kainui):

Not right at all unfortunately. Have you heard of the power rule? It basically just states that when you take a derivative you take the exponent and multiply it out front and then subtract the exponent by 1. So for instance the derivative of x^3 is 3x^2. Similarly the derivative of 8/x is the same as 8*x^(-1) so the derivative is actually -8x^(-2) so the answer is -8/(9^2).

OpenStudy (anonymous):

Okay, so then how do we do. The position of an object at time t is given by s(t) = 1 - 12t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (sleepyhead314):

power rule like like x^n = n*x^(n-1) so x^1 = 1 and -12x = -12

OpenStudy (anonymous):

Ookay...? What else..

OpenStudy (sleepyhead314):

the derivative of any number without a variable is always 0 so derivative of 5 is 0 derivative of -1243425 is 0 and derivative of 1 is 0

OpenStudy (anonymous):

so 12 disappears?

OpenStudy (sleepyhead314):

the "1" part of it would disappear the "-12x" part turns into -12 from the idea I described before remember?

OpenStudy (anonymous):

Right... So, then what do I after that? Do I plug in x?

OpenStudy (kainui):

\[1=x^0\] so a constant can be written as 12x^0. The power rule will just give you zero because the exponent is 0 when you multiply it out front.

OpenStudy (sleepyhead314):

well tell me what you think the derivative is...

OpenStudy (anonymous):

Okay, now I'm just confused. the derivative is -12, right?

OpenStudy (sleepyhead314):

yes! :D

OpenStudy (sleepyhead314):

they gave you a t=2 to trick you >,< all you needed was the derivative :P that'd be your answer ^_^

OpenStudy (anonymous):

Ok! Now, how do I find the velocity?

OpenStudy (kainui):

The derivative is the velocity, since velocity is the change in position.

OpenStudy (sleepyhead314):

lol xD velocity = derivative of position

OpenStudy (anonymous):

Oh! Alright then, thank you very much sleepyhead, have a nice day! c: